OpenStudy (anonymous):

Use the substitution method to solve the system of equations. Choose the correct ordered pair. 2x+4y=14 x=3

OpenStudy (anonymous):

Let see?

OpenStudy (anonymous):

plug in 2 for X and 5 for y so (2) + 2(5) = 12 2 + 10 = 12

OpenStudy (anonymous):

the second equation is already solved for x. So we can take this x=3 and substitute that in for every x in equation 1. 2(3)+4y=14. Now we can solve for y

OpenStudy (anonymous):

@Jerry123gnzcraft I don't know what Jerry is doing but it is not right

OpenStudy (anonymous):

2(3)+4y=14 6+4y=14 4y=8 y=2 (3,2)

OpenStudy (anonymous):

i agree wit chme